If a 25-kW (rated) wind turbine has a three-phase, 480-V generator, what minimum size wire will be needed for each phase to connect the wind turbine to a load that is 50 m away? Remember, you need to count the length of wire down the 25-m tower. Peak power can be 30 kW. Calculate maximum current and reduce it by a factor of 1.7 because the system is three-phase. Each leg (wire) of the three-phase system carries 1/3 of the current. Use Table 7.1.
Answer #1

Total length of wire =50m + 25m =75 m

Generator is rated at 480 V, three-phase.
I = P/V = 30,000/480 = 63 A
Need to worry about peak values

Since it is a 3 phase ,we need 52A. ()

From table 7.1,we need size 6 wire. (it's a box corresponding to 50A and 76m.